OpenStudy (anonymous):

An equation for estimating the speed of a tsunami is, s= 3.1√d where d is the depth of the ocean in meters and s is the speed in meters per second. Find the depth of the ocean when the speed of the tsunami is 250 meters per second. Estimate to the nearest tenth.

OpenStudy (anonymous):

Given: \[\huge s= 3.1\sqrt{d }\] Now we need to find d,when s= 250 m/s so, \[\huge 250= 3.1\sqrt{d } \rightarrow \frac{250}{3.1} = \sqrt{d}\] Next squaring boht the sides

OpenStudy (anonymous):

well 250/3.1 = 80.64

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

\[\huge \rightarrow (\frac{250}{3.1})^2 = (\sqrt d)^2\] \[\huge \rightarrow \frac{62500}{9.61} = d\] \[\huge d= 6503.64\] Hence the depth of the ocean is 6503.64 metres @jadeedwardson